OpenStudy (anonymous):

The time intervel between the arrival of a trough of a tsunami (tide goes very far out) and the crest (a huge wave) is 15 minutes. If the tsunami is moving at 400 km/h, what is its wavelenght

OpenStudy (ash2326):

time betwwen crest and trough =15 minutes time between two crest is =30 minutes time period of wave=30 minutes or 1800 second frequency=1/time period =1/1800 hertz wave's speed=400 km/h or 111.11 m/s wavelength= speed/frequency=111.11/(1/1800)=200000 meters or 20 km

OpenStudy (anonymous):

thx!